I cant seem to figure out how to write this proportion for the word problem:The length of a rectangular garden is 3/2 its width. It takes 50 m of edging to create a border. Find the dimensions of the garden.
Thank you so much!! If know how to write the proportion I'm sure I can figureout how to get the answer. Thanks again!

Answer by stanbon(48568)

Let the width be "x"
Then the length is (3/2)x
Formula: Perimeter = 2(lenth)+2(width)
EQUATION:
2x+2(3/2)x=50
Divide through by "2" to get:
x + (3/2)x=25
(5/2)x=25
1/2 x=5
x=10 meters
3/2 x= 15 meters
Cheers,
Stan H.